OpenStudy (anonymous):

Is the sequence geometric? If so, identify the common ratio. 1/3

OpenStudy (anonymous):

1/3,2/9,4/27,8/21,16/243

OpenStudy (anonymous):

No it is not. Let's compare the sequences 1st = 1/3 2nd = 2/9 = (1/3)(2/3) 3rd = 4/27 = (1/3)(2/3)^2 4th = 8/21 = (1/3) (2/7) 5th = 16/243 = (1/3)(2/3)^4 If it were geometric, the third term would be 8/81 = (1/3) (2/3)^3

OpenStudy (anonymous):

2nd line, terms* not sequences